# Multi-label Scene Classification for Autonomous Vehicles: Acquiring and Accumulating Knowledge from Diverse Datasets

### v2

@@ VI-F Driving Scene Identification Examples
To supplement the ablation study, seven examples (a-g) are further provided, with their corresponding results detailed in TABLE IV. The baseline classification model employing Swin-B pretrained on ImageNet as its deep encoder demonstrates insufficient knowledge about driving scene identification. It fails in multi-label scene classification, only correctly identifying each scene with at most three attributes.
@@
The foundation model, which assimilates knowledge regarding driving scene identification from diverse single-label datasets, significantly outperforms the baseline model that lacks such knowledge, correctly identifying scenes with respect to four to six attributes. However, an observable limitation of the foundation model is its frequent misidentification of dry road surface conditions as snowy, daytime as dawn, non-work zones as work zones, other road types as local roads, and non-intersection or three-way intersections as four-way intersections. For instance, the foggy weather condition in scenario (d) may have led the foundation model to erroneously classify daytime as dawn and dry road surfaces as snowy. Furthermore, the vehicle’s close proximity to the intersection makes it difficult to differentiate three-leg and four-leg features. Scenario (f) depicts an interstate highway that is slightly more complex than typical interstate driving scenes. The advertisement banner on the left and the open construction space on the right likely lead the model to misidentify it as a four-way intersection on a local road. Additionally, the atypical work zone features and the relatively farther distance of the safety drums from the ego-vehicle fail the model in recognizing the work zone scenario. Overall, certain combinations of scene attribute labels are less prevalent in single-labeled datasets, thereby limiting the foundation model’s exposure and subsequent ability to accurately identify these specific scenes.
@@
Utilizing feature proximity to recommend new training samples that are similar to the test data, CAL adapts the foundation model quickly to multitasking, demonstrating superior ability in classifying driving scenes comprehensively, as illustrated in the two rightmost columns in TABLE IV. Notably, the model adapted via CAL with only two iterations corrects most misclassifications in the foundation model, demonstrating the effectiveness of CAL. Meanwhile, with iterations up to five, the classification model learns the more fine-grained feature distributions, thus improving its ability to identify driving scenes. For instance, in scenes a, b, d, and e, the foundation model misclassified the weather-related road condition as snowy. Assisted by CAL, now the adapted model can correctly recognize the category and reassign the class label. Furthermore, CAL mitigates dominant biases in feature distributions across single-label classification datasets. Specifically, in recognizing a four-way intersection, the model relies more on the complex behavior of traffic participants rather than on the actual geometric layout of the intersection (as seen in scenes c and g). CAL addresses this by providing additional class labels to supervise the learning of features that were previously either unseen or unaddressed in monotask learning.
@@ VI-G Comparison to SOTA Multitask Classification Models
In the end, the proposed KAA-CAL learning system was compared to SOTA methods. These models for comparison, which were reviewed in Section II-B, are now briefly summarized below.
TABLE V: Classification accuracy (%) comparison with SOTA methods on BDD100K and Honda HSD dataset
@@
  • ResNet18 [5]: A multitask model for scene classification that utilizes ResNet18 as a shared deep encoder was developed in [5]. It was applied to BDD100K for identifying Time of Day, Weather, and Scene Type.
  • ViT-B-16 [5]: ViT-B-16 is also used as a shared deep encoder for driving scene classification on BDD100K [5].
  • CF-NET [6]: CF-NET utilizes ResNetV2-50 as the shared encoder and incorporates one 256-unit fully connected layer for each classifier. Single-label classifiers augment the multi-label classifier through feature fusion and stacking.
  • REC-NET [4]: REC-NET employs EfficientNet-B2 as the shared encoder and lightweight task-specific classifiers that each consists of two fully connected layers.
To demonstrate the generalization of the KAA-CAL learning system, the comparative study was conducted on both BDD100K[10] and HSD[11]. BDD100K provides the opportunity for identifying three driving scene attributes, including Time of Day, Weather, and Scene Type. The HSD dataset includes class labels for four attributes: Intersection Type, Surface Condition, Driving Types, and Weather. Despite the limited number of attributes used for scene image labeling, these two datasets provide common benchmarks for objectively evaluating KAA-CAL against SOTA methods, whose performance is already reported. TABLE VI summarizes the sizes of these datasets in terms of training, validation, and testing partitions.
TABLE VI: BDD100K and HSD Datasets
Dataset Training Validation Testing
BDD100K 70,000 7,000 3,000
HSD 53,115 5,500 2,275
Scene images were sampled every 100 frames from 80 hours
  of driving videos.
Our multitask classification model utilizes the foundation model as its deep encoder. It was subsequently trained via CAL to identify driving scenes in BDD100K and HSD, respectively, using a total of 15% of their training data over three iterations. In contrast, the four SOTA models, whose encoders were pretrained on ImageNet, were trained on the full BDD100K and HSD training datasets, respectively. Results of the comparative study are summarized in TABLE V.
From TABLE V, it can be seen that KAA-CAL uses only 15% of the training data and achieves task accuracy comparable to SOTA methods that utilize all the training data. Specifically, on BDD100K, KAA-CAL achieves the same highest accuracy as ResNet18 and CF-NET in identifying Time of Day. KAA-CAL’s accuracy in identifying Weather (81.8%) exceeds the second-best method, CF-NET (81.0%), by 0.8%, and it slightly surpasses CF-NET by 0.1% in identifying Scene Type.
@@
Importantly, Scene Type in BDD 100K and Driving Type in HSD are not attributes annotated in DSI. Therefore, the foundation model was not trained to recognize them. With the foundation model’s comprehensive knowledge of driving scene identification, and facilitated by CAL, the classification model efficiently gains the knowledge to identify driving scenes using these new attributes.
@@ VII Conclusion
This paper introduced a new learning system called KAA-CAL, as well as the DSI dataset, to tackle multi-label driving scene classification, a challenging perception task for AVs. KAA acquires and accumulates knowledge about scene identification from the DSI dataset via monotask learning, improving the average classification accuracy by 31.3% compared to the baseline model pretrained on ImageNet. Subsequently, CAL resolves the domain shifts of feature and attribute distributions, boosting the performance to 92.7%. Additionally, achieving performance competitive with SOTA multitask models on BDD100K and HSD datasets while using only 15% of the training data highlights KAA-CAL’s generalization.
Several limitations are presented in this paper, which necessitate further investigation in future work. Primarily, due to resource constraints and a deliberate focus on methodological development, KAA-CAL was not evaluated on a broader range of datasets. Although AVs demand a more comprehensive approach to scene identification, this paper was dedicated to multi-label image classification, leaving perception tasks such as semantic segmentation and object detection to future exploration. Furthermore, despite its efficiency, CAL has not yet achieved the ideal few-shot model refinement for downstream tasks. Nevertheless, the work presented in this paper builds a foundation for further exploring these opportunities for improvement. Future study can also leverage the domain-specific prior and semantic guidance embedded in the driving scene identification model to advance video understanding and reasoning, and motion prediction for AVs.
@@ Acknowledgment
Qin receives funding from the Rural Safe Efficient Advanced Transportation (R-SEAT) Center, a Tier-1 University Transportation Center (UTC) funded by the United States Department of Transportation (USDOT), through agreement number 69A3552348321. The contents of this paper reflect the views of the authors. USDOT assumes no liability for the contents or use thereof.

### v3

@@ VI-F Driving Scene Identification Examples
To supplement the ablation study in the preceding section, seven examples (a-g) are further provided, with their corresponding results detailed in TABLE IV. The baseline model demonstrates insufficient knowledge about driving scene identification. It fails in multi-label scene classification, only correctly identifying each scene with at most three attributes.
@@
The foundation model, which assimilates knowledge regarding driving scene identification from diverse single-label datasets, significantly outperforms the baseline model that lacks such knowledge, correctly identifying scenes with respect to four to six attributes. However, an observable limitation of the foundation model is its frequent misidentification of dry road surface conditions as snowy, daytime as dawn, non-work zones as work zones, other road types as local roads, and non-intersection or three-way intersections as four-way intersections. For instance, the foggy weather condition in scenario (d) may have led the foundation model to erroneously classify daytime as dawn and the dry road surface as snowy. Furthermore, the vehicle’s close proximity to the intersection makes it difficult to differentiate three-leg and four-leg features. Scenario (f) depicts an interstate highway that is slightly more complex than typical interstate driving scenes. The advertisement banner on the left and the open construction space on the right likely lead the model to misidentify it as a four-way intersection on a local road. Additionally, the atypical work zone features and the relatively farther distance of the safety drums from the ego-vehicle fail the model in recognizing the work zone scenario. Overall, certain combinations of scene attribute labels are less prevalent in single-labeled datasets, thereby limiting the foundation model’s exposure and subsequent ability to accurately identify these specific scenes.
@@
Utilizing feature proximity to recommend new training samples that are similar to the test data, CAL adapts the foundation model quickly to multi-attribute sence identification, demonstrating superior ability in classifying driving scenes comprehensively, as illustrated in the two rightmost columns in TABLE IV. Notably, the model adapted via CAL with only two iterations corrects most misclassifications by the foundation model, demonstrating the effectiveness of CAL. Meanwhile, with iterations up to five, the classification model learns the more fine-grained feature distributions, thus improving its ability to identify driving scenes. For instance, in scenes a, b, d, and e, the foundation model misclassified the weather-related road condition as snowy. Assisted by CAL, now the adapted model can correctly recognize the category and reassign the class label. Furthermore, CAL mitigates dominant biases in feature distributions across single-label classification datasets. Specifically, in recognizing a four-way intersection, the model relies more on the complex behavior of traffic participants rather than on the actual geometric layout of the intersection (as seen in scenes c and g). CAL addresses this by providing additional class labels to supervise the learning of features that were previously either unseen or unaddressed in monotask learning.
@@ VI-G Comparison to SOTA Multitask Classification Models
In the end, the proposed KAA-CAL learning method was compared to SOTA models reviewed in Section II-B. To demonstrate the generalization of the KAA-CAL learning method, the comparative study was conducted on both BDD100K [7] and HSD [8]. BDD100K has labels for three driving scene attributes, including Time of Day, Weather, and Scene Type. The HSD dataset includes labels for four attributes: Intersection Type, Surface Condition, Driving Types, and Weather. Despite the limited number of attributes provided, these two datasets offer an opportunity for objectively evaluating KAA-CAL against SOTA methods. TABLE V summarizes the sizes of these datasets in terms of training, validation, and testing partitions.
TABLE V: BDD100K and HSD Datasets
Dataset Training Validation Testing
BDD100K 70,000 7,000 3,000
HSD 53,115 5,500 2,275
Scene images were sampled every 100 frames from 80 hours
  of driving videos.
Our multi-attribute scene identification model utilizes the foundation model as its deep encoder. It was subsequently refined via CAL to identify driving scenes in BDD100K and HSD, respectively, using a total of 15% of their training data over three iterations. In contrast, the four SOTA models, whose encoders were pretrained on ImageNet, were trained on the full BDD100K and HSD training datasets, respectively. Results of the comparative study are summarized in TABLE VI.
TABLE VI: Classification accuracy (%) comparison with SOTA methods on BDD100K and Honda HSD dataset
@@
The best performance is bolded, and the second best is underlined. $\ast$ trained using only 15% of the training data.
KAA-CAL, which uses only 15% of the training data, achieves task accuracy comparable to SOTA methods that utilize all the training data. Specifically, on BDD100K, KAA-CAL achieves the same highest accuracy as ResNet-18 and CF-NET in identifying Time of Day. KAA-CAL’s accuracy in identifying Weather (81.8%) exceeds the second-best method, CF-NET (81.0%), by 0.8%, and it slightly surpasses CF-NET by 0.1% in identifying Scene Type.
@@
Notably, Scene Type in BDD 100K and Driving Type in HSD are not attributes that the foundation learned to recognize. With the foundation model’s comprehensive knowledge about driving scenes, and facilitated by CAL, the model efficiently gains the knowledge to identify driving scenes using these new attributes. While SOTA models listed in TABLE VI attain the performance comparable to KAA-CAL, they are not ready to characterize driving scenes using other attributes.
@@ VII Conclusion
This paper presents KAA-CAL, a deep learning method for multi-attribute driving scene identification, a fundamental yet challenging visual perception capability for AVs. KAA improved the average classification accuracy by 31.3% compared to the baseline model pretrained on ImageNet. CAL further boosts the performance to 92.7%. Additionally, KAA-CAL outperforms SOTA methods on BDD100K and HSD, achieving this while using only 15% of the training data and even recognizing attributes for which the foundation models was not previously train.
KAA-CAL lays a methodological foundation for not only multi-attribute scene identification - a high level perception - but other vision capabilities that AV needs. A universal vision foundation model can be built by synergizing the knowledge of those constituents. Due to a deliberate focus on the methodological development, KAA-CAL was not evaluated on a broader range of datasets. Yet, the number of scene attributes can be scaled up in this foundation model by following the data preparation and learning approach outlined in this paper. Furthermore, despite its efficiency, CAL has not yet achieved the ideal few-shot model refinement for downstream tasks. Future studies can also leverage the domain-specific prior and semantic guidance embedded in the driving scene identification model with vision-language models to advance video understanding and reasoning for AVs. Nevertheless, the work presented in this paper lays the groundwork for exploring these opportunities for generalization, scaling, improvement, and extension.
@@ Acknowledgment
Qin receives funding from the Rural Safe Efficient Advanced Transportation Center, a Tier-1 University Transportation Center funded by the United States Department of Transportation (USDOT), through agreement number 69A3552348321. The contents of this paper reflect the views of the authors. USDOT assumes no liability for the contents or use thereof.